(3,-6) x+y=-4
Write equation of the line that passes thru the given point and is perpendicular to the graph

Answer by stanbon(75887)

(3,-6) x+y=-4
Write equation of the line that passes thru the given point and is perpendicular to the graph
--------------
Given Equation::
x + y = -4
y = -x -4
slope = -1
------
xDesired equation::
slope = 1
Form: y = mx + b
-6 = 1*3 + b
b = -9
--------
Equation:
y = x - 9
--------------
Cheers,
Stan H.
-------------------